OpenStudy (anonymous):

Evaluate the expression: 3^x =26 which is true? a) 2 < x < 3 b)2 < x c) 3 < x < 4 d) 5 < x

OpenStudy (bahrom7893):

3^2 = 9 3^3 = 27

OpenStudy (bahrom7893):

does that show you anything?

OpenStudy (anonymous):

3^2 = 9 < 26 and 3^3 = 27 > 26, so 2 < x < 3. Obviously A. Hope that helped!

OpenStudy (anonymous):

Thanks